OpenStudy (anonymous):

In how many ways can the letters of the word 'principle' be arranged? In what proportion of these arrangements do the letters 'p' come together?

OpenStudy (nurali):

Permulation Formula \[nPr=\frac{ n! }{ (n-r)! }\] There are 9 letters in the word PRINCIPLE. 9P9=90720

OpenStudy (cwrw238):

if every letter was different there would be 9! ways but there are 2p's and 2 i's in principal - do you know how to cater for this?

OpenStudy (uri):

PRINCIPLE=9 So 9!/2!2! Because there are 2 p's and 2 i's

OpenStudy (anonymous):

I did that, its wrong.. the answer is 90, 720; 2/9

OpenStudy (uri):

The answer is 90,720.

OpenStudy (anonymous):

my bad sorry i read the answer wrong, thankyou x

OpenStudy (uri):

9*8*7*6*5*4*3*2*1=362880 and 2!2!=2*1 2*1=4 362880/4=90,720.